OpenStudy (anonymous):

integral dx over square root of quantity of 2x-x^2

OpenStudy (anonymous):

Focus on the quantity within the square root You can rewrite 2x -x^2 as 1- (x^2 -2x +1) this gives you \[\int\limits_{}^{}dx / (\sqrt{1-(x-1)^{2}}\] now put x-1=sin t dx=cost dt the above expression converts to costdt/[cost] =\[\int\limits_{}^{}\]dt =t =\[\sin^{-1} (x-1)\]

OpenStudy (anonymous):

you are my hero!!!